1. Find the volume of the parallelpiped spanned by the vectors
A=i+j-k; B= i-j+k; C=-i+j+k.
1
Expert's answer
2020-09-15T16:08:18-0400

The volume of a parallelepiped determined by the vectors a, b, and c, is the magnitude of their scalar triple product.

and the formula for finding the volume is given by;

Volume = (height) × (Area of base)

"Volume, V=\begin{vmatrix}
 𝐚. (𝐛 ×𝐜) \\
 
\end{vmatrix}"


"⇒V=\begin{vmatrix}
 (1,1,-1).((1,-1,1)×(-1,1,1)) \\
 
\end{vmatrix}"

The scalar triple product can be calculated using the formula;

"𝐚.(𝐛 × 𝐜) = \begin{vmatrix}
 ai & aj & ak \\
 bi & bj & bk \\
ci & cj & ck
\end{vmatrix}"


"=\begin{vmatrix}
 1 & 1 & -1 \\
 1 & -1 & 1 \\
-1 & 1 & 1
\end{vmatrix}"

"= -1(1\u00d71-(-1)\u00d7(-1))-1(1\u00d71-(-1)\u00d71)+1(1\u00d7(-1)-1\u00d71)"

"=-1(1-1)-1(1+1)+1(-1-1)"

"=-1(0)-1(2)+1(-2)"

"=0-2-2"

"= -4"

Since we want the modulus, i.e. the absolute value of the result, we therefore conclude that the parallelpiped has volume 4 cubic units.